# Modeling dissipation in quantum active matter

### v1

@@ II.2.1 Static Lindblad dissipator
with dissipation strength then regulated as
$\gamma=2(\nu_{-}-\nu_{+}).$ (9b)
The dissipator (II.2.1) can then also be written as
$\hat{\mathcal{D}}_{L}$ $=$ $\nu_{+}\left(\hat{a}^{\dagger}\hat{\rho}(t)\hat{a}-\frac{1}{2}\left\{\hat{a}\hat{a}^{\dagger},\hat{\rho}(t)\right\}\right)$ (10)
$+$ $\nu_{-}\left(\hat{a}\hat{\rho}(t)\hat{a}^{\dagger}-\frac{1}{2}\left\{\hat{a}^{\dagger}\hat{a},\hat{\rho}(t)\right\}\right).$
The idea of [5] was to control the quantum particle using a trapping potential to mimic active motion at long timescales, while focusing on possible quantum effects that emerge at short times, when the particle remains localized near $x=0$ and the creation-annihilation operators act on a particle at that position. The static Lindblad master equation (II.2.1) is suitable for this purpose under weak dissipation, since (i) the dissipative contribution is most relevant in the short-time regime, when the particle remains close to $x=0$ and the dissipator significantly affects the dynamics, and (ii) at longer times, the weak coupling to the bath implies that the effect of the dissipator is dominated by Hamiltonian dynamics.
However, as the dissipation increases, the contribution proportional to $\nu_{-}$ in the dissipator $\hat{\mathcal{D}}$ becomes dominant. This term, associated with the annihilation operator, drives the particle toward the ground state at $x=0$ rather then to minimum of the driving potential $x=x_{c}$, thus reducing the effectiveness of the trapping potential in controlling the quantum particle.
II.2.2 Translated Lindblad master equation
One possible extension of the framework is to modify the dissipator (II.2.1) in such a way that the new quantum master equation has its instantaneous steady-state as a Boltzmann state with the harmonic potential centered at $x_{c}(t)$. This modification would allow the dynamics to properly account for the trap motion on long timescales. For this purpose, we consider a dissipator $\hat{\mathcal{D}}$ translated by $x_{c}$. This translation is mediated by a translation operator $\hat{T}(x_{c})=e^{-ix_{c}\hat{p}/\hbar}$, the application of which yields (see Appendix A for details of the analytical calculations)
$\hat{\mathcal{D}}_{TL}$ $\equiv$ $\hat{\mathcal{D}}(x_{c}(t);\hat{\rho}(t))=\hat{T}(x_{c}(t))\hat{\mathcal{D}}(\hat{\rho}(t))\hat{T}^{\dagger}(x_{c}(t))$
$=$ $\frac{\gamma}{2}\bar{n}\left(\hat{\tilde{a}}^{\dagger}(t)\hat{\rho}(t)\hat{\tilde{a}}(t)-\frac{1}{2}\left\{\hat{\tilde{a}}(t)\hat{\tilde{a}}^{\dagger}(t),\hat{\rho}(t)\right\}\right)$
$+$ $\frac{\gamma}{2}(\bar{n}+1)\left(\hat{\tilde{a}}(t)\hat{\rho}(t)\hat{\tilde{a}}^{\dagger}(t)-\frac{1}{2}\left\{\hat{\tilde{a}}^{\dagger}(t)\hat{\tilde{a}}(t),\hat{\rho}(t)\right\}\right).$
Here, the modified creation and annihilation operators $\hat{\tilde{a}}^{\dagger}(t),\hat{\tilde{a}}(t)$ are adjusted to the position of the potential minimum $x_{c}(t)$
@@ II.2.2 Translated Lindblad master equation
We note that for the mathematical consistency of the suggested approach, $\dot{x}_{c}(t)$ must be a continuous function so that the driving protocol is continuous and differentiable in time (the necessity of this condition is discussed in Appendix C). The process $x_{c}(t)$ set by Eq. (1) satisfies this condition.
II.2.3 Agarwal dissipator
As an alternative to Lindblad dynamics, we consider the Agarwal dissipator [40, 41, 42, 43], which in the high-temperature limit recovers the well-known Caldeira-Leggett model [29]. While not strictly CPTP and thus not positivity-preserving, the Agarwal dissipator allows (in contrast to the Lindblad ones) to reproduce the classical Fokker-Planck equation in the classical limit. Moreover, it yields the correct thermodynamics. This makes it well-suited for studying classical systems approaching regimes where quantum effects become significant. In contrast, the Lindblad master equation with the CPTP density matrix is appropriate for genuinely quantum setups – such as quantum optics experiments – but fails to give the correct thermodynamics for non-Gaussian states or non-harmonic Hamiltonians.
The dissipator $\hat{\mathcal{D}}$ for the Agarwal dissipator reads:
$\hat{\mathcal{D}}_{A}$ $\equiv$ $\hat{\mathcal{D}}(\hat{\rho}(t))$
We note that the translational operator $\hat{T}$ does not affect the dissipator (II.2.3) (see Appendix A), and thus the dissipator (II.2.3) is translation invariant.
II.3 Numerical procedure
We choose $\sqrt{\hbar/m\omega}$ and $\omega^{-1}$ as units of length and time, respectively. In what follows, we compare the numerical results for the quantum mean squared displacements (MSDs), defined as [5]

### v2

@@ II.2.1 Lindblad dissipator
with thermalization rate then regulated as
$\gamma=2(\nu_{-}-\nu_{+}),$ (9b)
and temperature as
$T=\frac{\hbar\omega}{k_{B}\ln\left(\frac{\nu_{-}}{\nu_{+}}\right)}$ (9c)
The dissipator (II.2.1) can be conveniently rewritten as
${\mathcal{D}}_{L}\!$ $=$ $\!\nu_{+}\!\left(\hat{\tilde{a}}^{\dagger}(t)\hat{\rho}(t)\hat{\tilde{a}}(t)-\frac{1}{2}\left\{\hat{\tilde{a}}(t)\hat{\tilde{a}}^{\dagger}(t),\hat{\rho}(t)\right\}\right)$ (10)
$+$ $\!\nu_{-}\!\left(\hat{\tilde{a}}(t)\hat{\rho}(t)\hat{\tilde{a}}^{\dagger}(t)-\frac{1}{2}\left\{\hat{\tilde{a}}^{\dagger}(t)\hat{\tilde{a}}(t),\hat{\rho}(t)\right\}\right)\!.$
A convenient way to visualize the dynamics that dissipator (II.2.1) generates makes use of the Wigner quasi-probability function $W(x,p)$, which provides a convenient phase-space representation of the state of a quantum particle [46, 45, 47].
In the Wigner representation, the quantum master equation (II.1) corresponds to a Fokker-Planck equation for $W(x,p)$, providing an alternative but fully equivalent description of the same quantum dynamics (see Appendix A.1 for the derivation). The Fokker-Planck equation then reads:
$\partial_{t}W\!$ $=$ $\!\sum_{i=x,p}\partial_{i}(f_{i}W)+\sum_{i,j=x,p}\partial^{2}_{ij}(g_{ij}W),$ (11a)
$f_{x}\!$ $=$ $\!-\frac{p}{m}\!+\!\frac{\gamma}{4}(x-x_{c}),\,f_{p}\!=\!m\omega^{2}(x-x_{c})\!+\!\frac{\gamma}{4}p$ (11b)
$g_{xx}\!$ $=$ $\!\frac{\hbar\gamma}{8m\omega}\coth\!\left(\!\frac{\hbar\omega}{2k_{B}T}\!\right),\;g_{pp}\!=\!\frac{\hbar\gamma m\omega}{8}\coth\!\left(\!\frac{\hbar\omega}{2k_{B}T}\!\right),$
while $g_{xp}=g_{px}=0$. We note that the Wigner Fokker-Planck equation contains terms that are not present in a dissipative classical harmonic oscillator (CHO),
$m\ddot{x}_{\rm cho}(t)$ $+$ $\gamma\dot{x}_{\rm cho}(t)$
$=$ $-m\omega^{2}(x_{\rm cho}(t)-x_{c})+\sqrt{\gamma k_{B}T}\eta(t)\,$,
for which the Fokker-Planck equation reads [48]:
$\partial_{t}P_{\rm cho}\!$ $=$ $\!\sum_{i=x,p}\partial_{i}(f_{i}P_{\rm cho})+\!\sum_{i,j=x,p}\partial^{2}_{ij}(g_{ij}P_{\rm cho}),$ (13a)
$f_{x}\!$ $=$ $\!-\frac{p}{m},f_{p}=m\omega^{2}(x-x_{c})+\gamma p,\,$ (13b)
$g_{pp}\!$ $=$ $\!\gamma k_{B}T,\,g_{xx}=g_{xp}=g_{px}=0.\;$ (13c)
To make the extremum points of the steady-state probability distribution explicit, here we take that $x_{c}$ is constant so that the probability distribution of the quantum particle’s positions is centered at $x_{c}$ in the steady state. For fixed $x_{c}$, Eq. (11) has the stationary solution
$W(x,p)\!$
$=$ $\mathcal{N}\exp\left[-\tanh\left(\!\frac{\hbar\omega}{2k_{B}T}\!\right)\!\left(\frac{m\omega(x\!-\!x_{c})^{2}}{\hbar}\!+\!\frac{p^{2}}{m\hbar\omega}\right)\right]\,.$
The steady-state Wigner function, shown in Fig. 2 in this case has the extremum of the distribution at $x=x_{c}$ and $p=0$ for both weak (Fig. 2(a)) and strong (Fig. 2(b)) dissipation strengths. Note that in our numerical studies $x_{c}=x_{c}(t)$ evolves with time; consequently, the particle is constantly driven away from the instantaneous trap center and is exposed to the colored noise mediated by the trap’s motion.
Figure 2: Steady-state Wigner function $W(x,p)$ for Lindblad master equation (II.2.1). The position of the trapping potential is fixed at $x_{c}=3$ for (a) weak dissipation and (b) strong dissipation. The amplitude of the Wigner function values is given by the corresponding color representation; the graphs on the left and right correspond to the peak of the Wigner function at the extremum in $x$ and $p$, respectively. The green dashed lines indicate the position ($x=x_{c}$) of the fixed ($p=0$) harmonic potential. The thermalization rate $\gamma=2(\nu_{-}-\nu_{+})$ (9b) is varied at fixed temperature $T=\hbar\omega/\left(k_{B}\ln\left[\nu_{-}/\nu_{+}\right]\right)$ (9c) with $\nu_{-}/\nu_{+}=100$ for $\nu_{+}=10^{-4}$ (weak) and $\nu_{+}=10^{-1}$ (strong).
Although the Lindblad master equation with dissipator (II.2.1) correctly reproduces active motion even under strong dissipation, it does not display classical active motion in the classical limit. In fact, the structure of (11) is different from that of (13) in the limit $\hbar\to 0$, specifically in the force fields $f_{i}$.
II.2.2 Agarwal dissipator
As an alternative to Lindblad dynamics, we consider what we denote by “Agarwal dissipator” [43, 30, 49, 50]:
${\mathcal{D}}_{A}$ $\equiv$ ${\mathcal{D}}(\hat{\rho}(t))$
This dissipator recovers the well-known Caldeira-Leggett model in the high-temperature limit [44]. We note that dissipator (II.2.2) is translationally invariant (see Appendix B).
The corresponding Wigner function representation reads (see Appendix A.3 for the derivation):
$\partial_{t}P\!$ $=$ $\!\sum_{i=x,p}\partial_{i}(f_{i}P)+\sum_{i,j=x,p}\partial^{2}_{ij}(g_{ij}P),$ (16a)
$f_{x}\!$ $=$ $\!-\frac{p}{m},f_{p}=m\omega^{2}(x-x_{c})+\gamma p\,$ (16b)
$g_{pp}\!$ $=$ $\!\frac{\hbar\gamma m\omega}{8}\coth\left(\frac{\hbar\omega}{2k_{B}T}\right),g_{xx}=g_{xp}=g_{px}=0.$
which is analogous to the classical dissipative harmonic oscillator (13), with both the force fields $f_{i}$ and diffusion coefficient $g_{ij}$ identical between the classical and quantum cases in the limit $\hbar\to 0$.
The stationary solution of Eq. (16) for fixed $x_{c}$ is given by Eq. (II.2.1) and therefore identical to that of the Lindblad master equation. However, while both Fokker-Planck equations have the same steady-state for $x_{c}$ constant, their dynamics differ. This leads to further discrepancies under a time-dependent driving protocol $x_{c}(t)$. In fact, in the dynamics governed by the Lindblad master equation with dissipator (II.2.1), both the Hamiltonian part and the dissipator $\mathcal{D}$} drive the quantum particle toward the minimum of the moving trap $x_{c}(t)$. In contrast, for the Agarwal dissipator, the dissipator $\mathcal{D}$} acts purely as a source of friction, as is evident from the corresponding Wigner function representations (11), (16).
Thus, even though the dissipator (II.2.2) does not satisfy the Lindblad theorem, it yields the correct thermodynamics. This makes it well-suited for studying classical systems approaching thermodynamic equilibrium where quantum effects become significant.
II.3 Numerical procedure
In order to assess quantum active behavior in our model, we compare the numerical results for the quantum mean squared displacements (MSDs), defined as [5]
[…]
where $\overline{\cdots}$ denotes the average over prescribed trajectories $x_{c}(t)$. In (17), we first evaluate the trace of the observable (i.e., perform the quantum averaging for given instantaneous $x_{c}(t)$), and only afterwards average over the prescribed trajectories $x_{c}(t)$. This ordering is natural since $x_{c}(t)$ is a classical variable, following a classical probability distribution. Therefore, Eq. (4) involves purely classical sampling, and the order of averaging in Eq. (17) is irrelevant.
[…]
II.4 Potential experimental realization
A possible realization of the model considered here is a quantum-mechanical particle in a trap centered at position $x_{c}(t)$. The trajectory of the center $\{x_{c}(t)\}$ can be generated by a computer. By repeating this procedure for a set of $\{x_{c}(t)\}$, the average value of the quantum mean squared displacement, explained in the following subsection, can be measured. A sketch of such experimental realization is shown in Fig. 1. We note that equations similar to the dissipator Eq. (II.2.1) are used to describe the laser cooling dynamics of a particle trapped by a harmonic oscillator, such as for instance a single trapped ion [51]. Here, the thermalization rate and the temperature can be varied by changing the intensity and detuning of the laser [52, 51, 53].
III Results for the mean-squared displacements
In this section, we present the time evolution of the MSD, Eq. (17) by comparing the predictions of both forms of dissipators considered in this paper. We analyze different thermalization rates $\gamma$, which we refer to as dissipation strengths by analogy with classical mechanics, since $\gamma$ plays the role of dissipation in the corresponding classical system (13). We choose $\sqrt{\hbar/m\omega}$ and $\omega^{-1}$ as units of length and time, respectively. In Fig. 3, we first compare the MSD of Eq. (17) of the master equation with Lindblad dissipator (II.2.1), using the classical MSD of the driving protocol, $\overline{x_{c}^{2}(t)}$, as a reference. We recall that $x_{c}(t)$ in Eq. (1) corresponds to a passive inertial particle, exhibiting $t^{3}$-scaling at short times before crossing over to diffusion [54]. When $x_{c}(t)$ acts as a colored-noise input to the moving potential, it drives the particle in a way that breaks the detailed balance, allowing us to study how quantum dissipators influence motion that mimics active dynamics.
For weak thermalization rates, $\gamma\ll 1$ the MSD initially shows diffusive behavior, followed by a crossover to parabolic growth and later back to linear diffusion, thereby exhibiting properties of an active particle [11] and consistent with the predictions for low temperatures and weak thermalization [5]. However, as $\gamma$ increases, the effect of the dissipator $\mathcal{D}$} becomes more significant for $\gamma\gg 1$, so that the initial diffusive regime is no longer observable. Nevertheless, the ballistic-to-linear scaling of the MSD for $t\gtrsim\tau$ remains evident across the entire range of dissipation rates.
Figure 3: MSD of quantum-active particle for Lindblad dissipator (II.2.1) for weak, intermediate and strong dissipation respectively. The parameters of $x_{c}$ are $D=0.01,\tau=10$, and the corresponding classical $\operatorname*{MSD}=\overline{x_{c}^{2}(t)}$ of these trajectories is shown as a black line. The thermalization rate $\gamma=2(\nu_{-}-\nu_{+})$ (9b) is varied at fixed temperature $T=\hbar\omega/\left(k_{B}\ln\left[\nu_{-}/\nu_{+}\right]\right)$ (9c) with $\nu_{-}/\nu_{+}=100$ for $\nu_{+}=10^{-4}$ (weak), $\nu_{+}=10^{-2}$ (intermediate) and $\nu_{+}=10^{-1}$ (strong).
Figure 4: MSD of quantum-active particle with Agarwal dissipator (II.2.2) for weak, intermediate and strong dissipation respectively. The parameters of $x_{c}$ are $D=0.01,\tau=10$, and the corresponding classical $\operatorname*{MSD}=\overline{x_{c}^{2}(t)}$ of these trajectories is shown as a black line. The thermalization rate $\gamma=2(\nu_{-}-\nu_{+})$ (9b) is varied at fixed temperature $T=\hbar\omega/\left(k_{B}\ln\left[\nu_{-}/\nu_{+}\right]\right)$ (9c) with $\nu_{-}/\nu_{+}=100$ for $\nu_{+}=10^{-4}$ (weak), $\nu_{+}=10^{-2}$ (intermediate) and $\nu_{+}=10^{-1}$ (strong).
The differences between the predictions of the Lindblad and of the Agarwal dissipators are visible by comparing Fig. 3 and Fig. 4 and become evident at the dissipation time scale of the classical particle $\tau$. At long time scales, $t\gtrsim\tau$, similar to the Lindblad dissipator, the dynamics of quantum particles also follows that of the center of the trapping potential and exhibits the features of the active motion across all dissipation rates. However, the different nature of the considered dissipators leads to distinct MSD regimes at short time scales, $t\ll\tau$. In the Lindblad case, quantum fluctuations induced by the dissipator – generated by the creation and annihilation operators of the Lindblad dissipator – produce a diffusive regime (Fig. 3). By contrast, this short-time diffusive regime is absent for the Agarwal dissipator (Fig. 4): the quantum particle simply follows the stochastic trajectory $x_{c}(t)$ with an inertial delay and has the same $t^{3}$-scaling as the driving protocol.
IV Conclusions and outlook
In this work, we studied how different types of quantum dissipators influence the emergence of active-like dynamics of a quantum particle, aiming to understand how the environment affects the characteristics of motion reminiscent of classical active matter. We have compared the dynamics predicted by three forms of time-local master equation and explored the interplay between dissipation strength, external driving, and quantum effects on determining the mean squared deviation. The dissipators individually satisfy different requirements: either positivity or the correct thermodynamic properties. We find that the quantum particle exhibits typical behaviors of active particles, namely ballistic motion at intermediate times ($\textrm{MSD}\propto t^{2}$) and active diffusion at late times ($\textrm{MSD}\propto t$). Importantly, after averaging over many realizations of the actuating potential, the system is isotropic, such that the external field does not break the symmetry of the system, which is typical for active matter systems [8]. Beyond that, the short-time dynamics is governed by quantum effects and depends, in particular, on the chosen form of the dissipator.
Our findings could be experimentally tested in cold-atom systems, in particular in moving optical traps with controlled fluctuations of the laser intensity [55]. From a theoretical perspective, developing strategies for controlling quantum particles in stochastic [56] and dissipative [45] environments will be crucial for bridging the gap between abstract models and experimentally accessible setups.
Acknowledgments
M.t.V. is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – SFB 1551, Project-ID 464588647. G.M. acknowledges funding from the QuantERA II Programme (project “QNet: Quantum transport, metastability, and neuromorphic applications in Quantum Networks”), which has received funding from the EU’s Horizon 2020 research and innovation programme under Grant Agreement No. 101017733, as well as from the Deutsche Forschungsgemeinschaft DFG (Project ID 532771420). The authors acknowledge discussions with L. Chomaz, T. Pfau, P. Schmelcher, and A. Widera. S.L. thanks R. Kosloff for a discussion on modeling an open quantum system.
Appendix A Wigner-Weyl transform
A.1 Lindblad dissipator
We consider here the creation-annihilation operators in the form
@@ A.1 Lindblad dissipator
Here, we introduced a parameter $\delta$ in order to cover both the scenario studied in the main text ($\delta=1$, cf. Eq. (6)) and the scenario studied in Ref. [5] ($\delta=0$), see Appendix A.2 for a discussion how these forms are related.
$\left(\partial_{t}\rho\right)_{W}$ $\equiv$ $\partial_{t}W(x,p,t),$ (19a)
where $\star$ is the Moyal product defined as:
or equivalently, expanded in powers of $\hbar$:
${\frac{1}{i\hbar}}(f\star g-g\star f)=\{f,g\}+O(\hbar^{2}).$ (22)
For the Lindblad dissipator,
$\left(\hat{A}\hat{\rho}\hat{B}\right)_{W}$ $=$ $(A_{W}\star W)\star B_{W}$ (23)
Given that operators $\hat{A},\hat{B}$ in the dissipator $\mathcal{D}$} are linear in $x,p$, Eq. (24) reduces to:
$\left(\hat{A}\hat{\rho}\hat{B}\right)_{W}$ $=$ $A_{W}WB_{W}$ (25)
Thus, the triple products relevant in our case are
$\left(\hat{\tilde{a}}^{\dagger}\hat{\rho}\hat{\tilde{a}}\right)_{W}$ $=$ $|\tilde{a}|^{2}W$ (26)
$+$ $\frac{i\hbar}{2}\left((\alpha\beta-(-\beta)\alpha)W+\tilde{a}^{\dagger}(\partial_{x}W\beta-\partial_{p}W\alpha)+\tilde{a}(\alpha\partial_{p}W-(-\beta)\partial_{x}W)\right)$
$=$ $|\tilde{a}|^{2}W-\frac{1}{2}\left(W+p\partial_{p}W+(x-\delta x_{c})\partial_{x}W\right)+\frac{m\omega\hbar}{8}\partial_{x}^{2}W+\frac{\hbar}{8m\omega}\partial_{p}^{2}W;$
$\left(\hat{\tilde{a}}\hat{\rho}\hat{\tilde{a}}^{\dagger}\right)_{W}$ $=$ $|\tilde{a}|^{2}W$ (27)
$+$ $\frac{i\hbar}{2}\left((-\alpha\beta-\beta\alpha)W+\tilde{a}(-\partial_{x}W\beta-\partial_{p}W\alpha)+\tilde{a}^{\dagger}(\alpha\partial_{p}W-\beta\partial_{x}W)\right)$
$=$ $|\tilde{a}|^{2}W+\frac{1}{2}\left(W+p\partial_{p}W+(x-\delta x_{c})\partial_{x}W\right)+\frac{m\omega\hbar}{8}\partial_{x}^{2}W+\frac{\hbar}{8m\omega}\partial_{p}^{2}W;$
For the anti-commutators, we obtain
$\frac{1}{2}\left\{\hat{\tilde{a}}\hat{\tilde{a}}^{\dagger},\hat{\rho}\right\}$ $=$ $\left(|\tilde{a}|^{2}+\frac{i\hbar}{2}(\partial_{x}\tilde{a}\partial_{p}\tilde{a}^{\dagger}-\partial_{p}\tilde{a}\partial_{x}\tilde{a}^{\dagger})\right)W-\frac{\hbar^{2}}{8}(\partial_{x}^{2}|\tilde{a}|^{2}\partial_{p}^{2}W-\partial_{p}^{2}|\tilde{a}|^{2}\partial^{2}_{x}W)$ (28)
$=$ $(|\tilde{a}|^{2}+\frac{1}{2})W-\frac{\hbar^{2}}{4}(\alpha^{2}\partial_{p}^{2}W-\beta^{2}\partial^{2}_{x}W),$
$\frac{1}{2}\left\{\hat{\tilde{a}}^{\dagger}\hat{\tilde{a}},\hat{\rho}\right\}$ $=$ $\left(|\tilde{a}|^{2}+\frac{i\hbar}{2}(\partial_{x}\tilde{a}^{\dagger}\partial_{p}\tilde{a}-\partial_{p}\tilde{a}^{\dagger}\partial_{x}\tilde{a})\right)W-\frac{\hbar^{2}}{8}(\partial_{x}^{2}|\tilde{a}|^{2}\partial_{p}^{2}W-\partial_{p}^{2}|\tilde{a}|^{2}\partial^{2}_{x}W)$ (29)
$=$ $(|\tilde{a}|^{2}-\frac{1}{2})W-\frac{\hbar^{2}}{4}(\alpha^{2}\partial_{p}^{2}W-\beta^{2}\partial^{2}_{x}W).$
Finally, we arrive at
$({\mathcal{D}})_{W}$ $=$ $\nu_{+}\Bigg(-\frac{1}{2}\left(2W+p\partial_{p}W+(x-\delta x_{c})\partial_{x}W\right)+\frac{m\omega\hbar}{4}\partial_{x}^{2}W+\frac{\hbar}{4m\omega}\partial_{p}^{2}W\Bigg)$ (30)
$\partial_{t}W\!$ $=$ $\!\sum_{i=x,p}\partial_{i}(f_{i}W)+\sum_{i,j=x,p}\partial^{2}_{ij}(g_{ij}W),$ (31a)
$f_{x}$ $=$ $-\frac{p}{m}+\frac{\gamma}{4}(x-\delta x_{c}),\quad f_{p}=m\omega^{2}(x-x_{c})+\frac{\gamma}{4}p$ (31b)
$g_{xx}$ $=$ $\frac{\hbar\gamma}{8m\omega}\coth\left(\frac{\hbar\omega}{2k_{B}T}\right),\quad g_{pp}=\frac{\hbar\gamma m\omega}{8}\coth\left(\frac{\hbar\omega}{2k_{B}T}\right),\quad g_{xp}=g_{px}=0.$
Here, we have used the following relations:
$\nu_{-}$ $=$ $\frac{\gamma}{2}(\bar{n}+1),\ \nu_{+}=\frac{\gamma}{2}\bar{n},$ (32)
A.2 Role of the Lindblad dissipator form
For the Lindblad dissipator, Ref. [5] has used instead of Eq. (11) the alternative form:
${\mathcal{D}}_{SL}={\mathcal{D}}(\hat{\rho}(t))$ $=$ $\frac{\gamma}{2}\bar{n}\left(\hat{a}^{\dagger}\hat{\rho}(t)\hat{a}-\frac{1}{2}\left\{\hat{a}\hat{a}^{\dagger},\hat{\rho}(t)\right\}\right)$
with the static ladder operators